OpenStudy (chris215):

What is the average rate of change of y with respect to x over the interval [-3, 5] for the function y = 2x + 2?

Directrix (directrix):

Look at the attachment for the formula.

Directrix (directrix):

If x is -3, then what is the value of f(-3) = 2*(-3) +2 ? Also, if x = -5, then what is the value of f(-5) = 2*(-5) + 2 ? Post what you get, okay?

OpenStudy (chris215):

f(-3)=-3 f(-5)=-8

OpenStudy (chris215):

i mean f(-3)=-4

jimthompson5910 (jim_thompson5910):

Rule: if f(x) is a linear function, then the average rate of change on any interval is going to be equal to the slope of the line.